### v1

# Adaptive Input Design for Nonlinear System Identification with Operational Constraints

Abstract
We consider the problem of joint input design and parameter estimation for identifying nonlinear system models through the sequential acquisition of measurements while adhering to system constraints. We utilize a receding horizon approach and propose a new scale-invariant input design criterion, which is tailored to continuously updated parameter estimates, along with a new sequential parameter estimator. We demonstrate the ability of the method to design informative experiments online, while steering the system within operational constraints.
keywords:
input design, adaptive design, constrained experiments
[…]
algorithmic thanks: TODO: Acknowledgments
1 Introduction
We consider nonlinear dynamical systems that can be modelled in discrete-time as
$x_{t+1}=f_{\theta}(x_{t},u_{t}),$ (1)
where both states $x_{t}$ and parameters $\theta$ are unknown. Using the inputs $u_{t}$ and measurements of the system outputs $y_{t}$, we wish to estimate $\theta$. The goal is to choose $u_{t}$ so as to make the data informative about $\theta$, while respecting the system’s operating constraints. This paper thus tackles the challenging joint problem of adaptive input design and parameter estimation for nonlinear systems.
Figure 1: Results using input designs for pendulum system (2) with $|u_{t}|\leq 10$ starting at sample $t=8$. (a) Example of angle $x_{t}^{(1)}$ using two different prbs inputs and an adaptive design (using L-criterion (8)) that steers the angle within the interval $[-45^{\circ},45^{\circ}]$. (b) Corresponding mean squared-errors of parameter estimator along with (approximate) Cramér-Rao bounds. The errors when using prbs2 are however too large to be included in the plot. Details are given in Section 4.1.
Input design for linear systems has been researched at least since the 1960s. A classic choice for $u_{t}$ is a pseudorandom binary sequence (prbs) [9] due to its excitation properties and ease of implementation. However, prbs signals do not utilize any information about the system. By leveraging this information more efficient inputs can be designed that reduce the experiment time. A statistically motivated approach uses the Fisher information matrix (fim) to design optimal input signals for linear systems [20, 28]. Since the fim, denoted $J$, determines a lower error bound for estimating $\theta$ via the Cramér-Rao inequality, input design methods typically use some scalar function of $J$ as design criterion for $u_{t}$. The most popular ones include the determinant $|J|$ and trace $\text{tr}\{J^{-1}\}$, referred to as D- and A-criteria, respectively [15, 6, 11]. A more tractable approach is to first design the input signals in the frequency domain, using the asymptotic error properties of prediction error methods, and then synthesize $u_{t}$ in the time domain [4, 12]. More recent frequency-domain approaches incorporate application-oriented criteria and constraints using advances in convex optimization techniques [5, 1].
The high cost of conducting identification experiments motivates the development of efficient input design schemes that continuously adapt $u_{t}$ to the information available up to time $t$. This implies a closed-loop method for joint design and identification. It was shown in [2, 14] that for finite-impulse response and autoregressive system models, optimal adaptive designs under amplitude constraints correspond to binary sequences [11]. More recently, an adaptive input design for output-error models was considered in [29]. Informative experiments may, however, have adverse effects on the closed-loop system behaviour. The balance between sufficiently perturbing input signals and control performance was initially addressed by Feldbaum [7] and is a trade off considered in recent work on nonlinear system identification [25, 3].
In this paper, our focus is on the adaptive design of $u_{t}$ that are informative about $\theta$ while steering the system states $x_{t}$ within operational constraints. This is particularly challenging in the case of nonlinear systems. For an illustration, consider a simple nonlinear model of a pendulum:
$\begin{split}x_{t+1}^{(1)}&=x_{t}^{(1)}+x_{t}^{(2)}\Delta{T}\\ x_{t+1}^{(2)}&=x_{t}^{(2)}+(\theta_{1}\sin{x_{t}^{(1)}}+\theta_{2}u_{t})\Delta{T},\end{split}$ (2)
where $x_{t}^{(1)}$ represents the angle from the vertical to the pendulum, $x_{t}^{(2)}$ denotes its angular velocity, and $\Delta{T}$ is a known sampling period. Suppose we wish to design an informative experiment that steers the pendulum angle within the interval $[-45^{\circ},45^{\circ}]$, despite both $\theta$ and $x_{t}$ being unknown. Figure 1 shows the resulting angles using a prbs (called prbs2) that is tuned to approximately satisfy the angular constraints during the experiment, and another prbs (called prbs1) that only satisfies input amplitude constraints. The former experiment is uninformative and yields unusable parameter estimates, whereas the latter is informative but grossly violates the specified state constraints.
Goodwin [10] introduced an open-loop input design method for nonlinear systems that penalizes designs in which the estimated states violate the constraints. This is achieved by linearizing the state-space model and designing $u_{t}$ over a fixed horizon. However, the method requires a good prior estimate of $\theta$ and the use of the A-criterion yields input designs that are not scale invariant with respect to $\theta$. More recently, Flila et al. [8] proposed the use of a receding-horizon approach to steer the nonlinear system within output constraints. This was achieved using a penalization technique and a design criterion that approximates $1/\text{tr}\{J\}$. The method was demonstrated for a thermal process model with temperature constraints. While computationally convenient, this design criterion is neither scale invariant nor does it lower bound the resulting estimation errors. Moreover, the statistical efficiency of the method is not studied.
In this paper, we aim to address the above challenges of adaptive designs for nonlinear system identification. Specifically, the principal contributions are:
  • a new scale-invariant design criterion tailored to the setting in which the parameter estimates are continuously updated;
  • an adaptive input design method with efficient parameter estimation that steers the system within operational constraints while achieving accurate estimates and reducing experiment time.
An illustration of the method was provided in Figure 1, where we see that the adaptive design steers the system states within operational constraint while yielding lower estimation errors than an unconstrained design.
The remainder of the paper is outlined as follows. First, we formalize the problem of interest and then turn to the proposed method. Statistical evaluations are presented using three different nonlinear systems, after which we conclude the paper.
2 Problem Formulation
We assume the following measurement model
$y_{t}=Hx_{t}+e_{t}$ (3)
where $y_{t}$ is a $d_{y}\times 1$ vector and $e_{t}$ is zero-mean white Gaussian process with a diagonal covariance matrix $V$. Sensor specifications typically provide accurate estimates of the noise variances. The dimensions of $u_{t}$, $\theta$ and $x_{t}$ are $d_{u}\times 1$, $d_{\theta}\times 1$ and $d_{x}\times 1$, respectively.
Let $\mathcal{D}_{1}^{t}=\{(u_{1},y_{1}),\dots,(u_{t},y_{t})\}$ denote a sequence of (possibly multivariate) input and output signals. The parameter vector $\theta$ in (1) is estimated using the likelihood function of $\mathcal{D}_{1}^{t}$.
An adaptive design of the input $u_{t+1}$ must take into account how the unknown states in (1) evolve in order to adhere to operational constraints and achieve efficient experiments. For a tractable formulation, we consider a receding horizon framework for designing a $k$-step sequence
$U_{t+1}=[u_{t+1},\dots,u_{t+k}].$ (4)
Starting from a state $x_{t+1}$, we then have a sequence
$X_{t+1}=[x_{t+1},\dots,x_{t+k+1}],$ (5)
when $U_{t+1}$ is applied to (1) for a given $\theta$.
We want both input and state sequences to adhere to certain constraints. Here we consider (element-wise) box constraints $u_{\min}\leq u_{t+i}\leq u_{\max}$ and $x_{\min}\leq x_{t+i}\leq x_{\max}$, which can be time varying. Note that the state constraints are nonlinear with respect to the inputs $U_{t+1}$. We denote the set of constrained inputs by $\mathcal{U}$ and constrained states by $\mathcal{X}$. We will assume that the constraints are feasible, viz. there exists an input sequence in $\mathcal{U}$ that yields a state sequence in $\mathcal{X}$.
A receding horizon formulation for input design is then
$U^{*}_{t+1}=\operatorname*{arg\,min}_{U_{t+1}\in\mathcal{U}}\mathcal{J}(U_{t+1};\theta,x_{t+1})\;\;\text{s.t. }X_{t+1}\in\mathcal{X},$ (6)
where $\mathcal{J}$ is some scalar function of the information matrix $J_{t+k}$ obtained from future samples
$\mathcal{D}^{t+k}_{t+1}=\{(u_{t+1},y_{t+1}),\dots,(u_{t+k},y_{t+k})\}.$
Since $\theta$ and $x_{t+1}$ are unknown we use plug-in estimates in (6) which turns $U^{*}_{t+1}$ into a design that is adapted to the data available up to time $t$.
In the next section, we formulate an appropriate design criterion and a corresponding relaxation of the constraints, that render a tractable adaptive input design method which steers the nonlinear system within operational constraints.
[…]
3 Method
We begin by formulating an adaptive design criterion and tractable relaxation of the state constraints in (6). We then turn to the parameter estimator and a way of approximating the solution to the receding-horizon design problem.
3.1 Design criterion and constraint relaxation
Suppose we have gathered data $\mathcal{D}^{t}_{1}$ and obtained an unbiased estimate $\widehat{\theta}_{t}$ with an error covariance $C_{t}$. With $k$ additional samples, we would have an updated estimate $\widehat{\theta}_{t+k}$ and we propose to use its relative error improvement as a design criterion. Specifically, the mean squared-error of $\widehat{\theta}_{t+k}$, standardized to the current errors, is bounded as follows:
$\mathcal{J}(U_{t+1};\theta,x_{t+1})\leq\mathbb{E}\left[\|\theta-\widehat{\theta}_{t+k}\|^{2}_{C^{-1}_{t}}\right]\leq d_{\theta},$ (7)
where
$\boxed{\mathcal{J}(U_{t+1};\theta,x_{t+1})=\text{tr}\left\{C^{-1}_{t}J^{-1}_{t+k}\right\},}$ (8)
which follows from the Cramér-Rao inequality [26, 19] and the fact that the error covariance matrix will be no worse than $C_{t}$. The weighted norm in (7) aligns the future errors to the axes of the current confidence ellipsoid. Note that the upper bound $d_{\theta}$ is attained in the extreme case of uninformative experiments. The criterion (8) is linear in $J^{-1}_{t+k}$ and convex in $J_{t+k}$, aka. an L-criterion [6, ch. 2.9] adapted to the current information.
We now turn to relaxing the state constraints, specified by $\mathcal{X}$ in (6). Instead of strict adherence, we instead aim for designs that continually steer the state to satisfy the operational constraints. This is achieved by introducing a penalty term to the design problem (6). Whereas $\mathcal{J}$ measures a standarized squared deviation of $\widehat{\theta}_{t+k}$ from the unknown $\theta$, we let penalty $\mathcal{J}_{\mathcal{X}}$ measure a standardized squared deviation of $x_{t+i}$ from the state boundaries. Specifically,
$\begin{split}&\mathcal{J}_{\mathcal{X}}(U_{t+1};\theta,x_{t+1})=\frac{1}{k}\times\\ &\sum_{i=1}^{k}\|(Sx_{\min}-Sx_{t+i})^{+}+(Sx_{t+i}-Sx_{\text{max}})^{+})\|^{2},\end{split}$ (9)
where $S$ is a diagonal matrix that standardizes deviations from the constraints. That is, $[S]_{i,i}=1/(x^{(i)}_{\max}-x^{(i)}_{\min})$ or 0, depending on whether $x_{i}$ is constrained. The relaxed form of (6) then becomes
$U^{*}_{t+1}=\operatorname*{arg\,min}_{U_{t+1}\in\mathcal{U}}\mathcal{J}(U_{t+1};\theta,x_{t+1})+\gamma\mathcal{J}_{\mathcal{X}}(U_{t+1};\theta,x_{t+1}),$ (10)
where $\gamma\geq 0$ is a gain that controls the adherence to the state constraints. Note that when the state sequence $X_{t+1}$ is in $\mathcal{X}$, the cost $\mathcal{J}_{\mathcal{X}}$ vanishes. By setting $\gamma\gg d_{\theta}$, the cost of deviating from the state boundaries quickly exceeds the cost of using uninformative inputs. This incentivizes designs that steer the system within the state constraint set $\mathcal{X}$.
To evaluate the criterion (8), we first turn to the computation of the fim. For notational compactness, we introduce the following variables:
$\beta_{t+1}=\begin{bmatrix}\theta\\ x_{t+1}\end{bmatrix}\quad\text{and}\quad\text{diag}(v)=V.$ (11)
Using the available data up to time $t$, we have corresponding estimates $\check{\beta}_{t+1}$ and $\widehat{v}_{t}$, respectively. We have used $\check{(\cdot)}$ to emphasize that $\check{\beta}_{t+1}$ contains a prediction of a future state (used in the estimator below).
Since the future $k$ samples are free variables and the corresponding measurements in $\mathcal{D}^{t+k}_{t+1}$ are independent of $\check{\beta}_{t+1}$ and $\widehat{v}_{t}$, we can express the log-likelihood of the data as:
$\ell(\mathcal{D}^{t+k}_{t+1},\check{\beta}_{t+1},\widehat{v}_{t}|\beta_{t+1},v)\\ =\ln p(y_{t+1:t+k}|U_{t+1},\beta_{t+1},v)+\ln p(\check{\beta}_{t+1},\widehat{v}_{t}|\beta_{t+1},v),$ (12)
where the last term summarizes the past information about the unknown $\beta_{t+1}$ and $v$. The log-likelihood yields a joint fim,
$J_{\beta_{t+1},v}=\mathbb{E}\left[\begin{bmatrix}\frac{\partial\ell}{\partial\beta_{t+1}}\\ \frac{\partial\ell}{\partial v}\\ \end{bmatrix}\begin{bmatrix}\frac{\partial\ell}{\partial\beta_{t+1}}\\ \frac{\partial\ell}{\partial v}\\ \end{bmatrix}^{\top}\right].$ (13)
We denote the error covariance matrices of $\check{\beta}_{t+1}$ and $\widehat{v}_{t}$ as $C_{\beta_{t+1}}$ and $C_{v_{t}}$, respectively, and their mutual error covariances are assumed to be negligible. Using Gaussian distributions [27, 23], (12) can be expressed as
$\ell=-\frac{1}{2}\|\beta_{t+1}-\check{\beta}_{t+1}\|_{C^{-1}_{\beta_{t+1}}}^{2}-\frac{1}{2}\|v-\widehat{v}_{t}\|^{2}_{C^{-1}_{v_{t}}}\\ -\frac{1}{2}\sum_{i=1}^{k}\|\varepsilon_{t+i}\|_{V^{-1}}^{2}-\frac{k}{2}\ln{|V|}+\text{const.}$ (14)
where $\varepsilon_{t}\equiv Hx_{t}-y_{t}$ is the prediction error. For the dynamic model (1), we define the composition of $i$ steps forward recursion by
$f^{i}_{\theta}(x_{t+1},u_{t+1:t+i})\\ =f_{\theta}(f_{\theta}(\dots f_{\theta}(x_{t+1},u_{t+1})\dots u_{t+i-1}),u_{t+i})$ (15)
Then using (1) and (3), we can express the prediction error as a function of $\beta_{t+1}$:
$\varepsilon_{t+i}$ $=Hx_{t+i}-y_{t+i}$
$=Hf^{i-1}_{\theta}(x_{t+1},u_{t+1:t+i-1})-y_{t+i}$
and define the Jacobian by
$E_{t+i}=H\frac{\partial f_{\theta}^{i-1}(x_{t+1},u_{t+1:t+i-1})}{\partial\beta_{t+1}}.$
Using (14) the joint fim (13) is partitioned into blocks as follows [26]
$\begin{split}J_{\beta_{t+1},v}&=\begin{bmatrix}J_{\theta}&J_{\theta x_{t+1}}&0\\ J_{\theta x_{t+1}}^{\top}&J_{x_{t+1}}&0\\ 0&0&J_{v}\end{bmatrix}\\ &=\begin{bmatrix}C_{\beta_{t+1}}^{-1}+\sum_{i=1}^{k}E^{\top}_{t+i}V^{-1}E_{t+i}&{0}\\ {0}&C_{v_{t}}^{-1}+\frac{k}{2}V^{-2}\end{bmatrix}\end{split}.$ (16)
This yields the sought information matrix for $\theta$ in (8):
$J_{t+k}\equiv J_{\theta}-J_{\theta x_{t+1}}J_{x_{t+1}}^{-1}J_{\theta x_{t+1}}^{\top},$ (17)
which is a function of the future inputs $U_{t+1}$ as well as the unknown $\theta$, $x_{t+1}$ and $v$. We note here that $x_{t+1}$, which initializes the state sequence, also affects the accuracy of estimating $\theta$. Next, we turn to the estimator of $\theta$ and its error covariance matrix $C_{t}$ used in (8).
3.2 Block-wise online estimator
The mle for the entire dataset $\mathcal{D}^{t}_{1}$ is intractable and the adaptive setting requires an online estimator. In the case of nonlinear systems, however, one-step online algorithms that jointly estimate states and parameters often suffer from poor initial convergence.
To alleviate this, we propose an online estimator of $\beta_{t}$ and $v$ that operates with $B$ sequential samples at a time while leveraging the prior estimate of the variables of interest at each sample block. We introduce the $b$th sample block as
$u^{b}=[u_{(b-1)B+1},\dots,u_{bB}]$
and $y^{b}$ is defined similarly, where $b=1,2,\dots$. Then $\beta^{b}\equiv\beta_{(b-1)B+1}$ includes the initial state $x_{(b-1)B+1}$ for block $b$. The log-likelihood for the $b$th data block is then similar to (12) and (14):
$\begin{split}&\ell(\mathcal{D}^{t}_{t-B+1},\check{\beta}^{b},\widehat{v}^{b-1}|\beta^{b},v)\\ =&\ln p(y^{b}|u^{b},\beta,v)+\ln{p(\check{\beta}^{b},\widehat{v}^{b-1}|\beta^{b},v)}\\ =&-\frac{1}{2}\|\beta^{b}-\check{\beta}^{b}\|_{\check{C}^{-1}_{\beta^{b}}}^{2}-\frac{1}{2}\|v-\widehat{v}^{b-1}\|^{2}_{C_{v^{b-1}}^{-1}}\\ &-\frac{1}{2}\sum_{i=0}^{B-1}\|\varepsilon_{t-i}\|_{V^{-1}}^{2}-\frac{B}{2}\ln{|V|}+\text{const.}\end{split}$ (18)
Note that $\check{\beta}^{b}$ contains a state prediction based on the data from the previous blocks, i.e., $\mathcal{D}_{1}^{t-B}$. This prediction is computed using the model (1) along with the unscented transform $T$ [13, 21], which is a statistical linearization technique that uses a deterministic sampling scheme based on the previous estimate $\widehat{\beta}^{b-1}$ and error covariance matrix $C_{\beta^{b-1}}$. We apply the transform $B$ forward steps in the model (1), that is,
$(\check{\beta}^{b},\check{C}_{\beta^{b}})=T^{B}(\widehat{\beta}^{b-1},C_{\beta^{b-1}};u^{b-1}),$ (19)
which provides $\check{\beta}^{b}$ along with an approximate error covariance matrix $\check{C}_{\beta^{b}}$.
Using the above quantities in (18) yields a block-wise estimator:
$(\widehat{\beta}^{b},\widehat{v}^{b})=\operatorname*{arg\,min}_{\beta^{b},v}\;-\ell(\mathcal{D}^{t}_{t-B+1},\check{\beta}^{b},\widehat{v}^{b-1}|\beta^{b},v)$ (20)
To minimize the non-convex cost function in (20) online, we used an efficient numerical search method that mitigates the impact of ‘rugged’ cost surfaces. Specifically, we applied the Basin Hopping algorithm [17, 22] to find an approximate solution of (20) using a quasi-Newton ‘limited-memory BFGS’ [18] minimizer for local searches.
Finally, since (20) approximates the maximum likelihood estimator, we use the fim based on (18) for the error covariance matrices, that is,
$C_{\beta_{t}}=\begin{bmatrix}J_{\theta}&J_{\theta x_{t}}\\ J_{\theta x_{t}}^{\top}&J_{x_{t}}\end{bmatrix}^{-1}\quad\text{and}\quad C_{v_{t}}=J^{-1}_{v},$ (21)
evaluated using plug-in estimates. Consequently, $C^{-1}_{t}=J_{\theta}-J_{\theta x_{t}}J_{x_{t}}^{-1}J_{\theta x_{t}}^{\top}$ is used in the adaptive L-criterion (8).
3.3 Adaptive design method
Using (20), we obtain estimates $\widehat{\theta}_{t}$, $\widehat{x}_{t+1}$, and $\widehat{v}_{t}$ which we plug in the cost functions in (10). We can turn the design problem into an unconstrained form by transforming the inputs $\mathcal{U}\rightarrow\mathbb{R}^{d_{u}}$ as follows:
$\widetilde{U}_{t+1}=[\widetilde{u}(u_{t+1}),\dots,\widetilde{u}(u_{t+k})]$ (22)
where
$\widetilde{u}(u_{t})=g\big{(}(u_{t}-u_{\text{min}})\oslash(u_{\text{max}}-u_{\text{min}})\big{)}$ (23)
and where $g$ is a differentiable and invertible function, e.g., the element-wise inverse sigmoid function $[g(z)]_{i}=\ln{\frac{z_{i}}{1-z_{i}}}$.
Thus (22) is unconstrained and computationally convenient for online input design. The map $\widetilde{U}_{t+1}$ recovers the original input space by:
$U_{t+1}=[\widetilde{u}^{-1}(\widetilde{u}_{t+1}),\dots,\widetilde{u}^{-1}(\widetilde{u}_{t+k})]$ (24)
where $\widetilde{u}^{-1}(\tilde{u}_{t})=g^{-1}(\tilde{u}_{t})\odot(u_{\max}-u_{\min})+u_{\min}$ and $[g^{-1}(z)]_{i}=\frac{1}{1+e^{-z_{i}}}$. Using the transformed inputs $\widetilde{U}_{t+1}$, by substituting (24) in (10), results in an unconstrained optimization problem. At $t$, the input
$u^{*}_{t+1}=\widetilde{u}^{-1}(u^{*}_{t+1})$
is computed and applied in the subsequent time step.
The online receding horizon problem (10) is tackled using a first-order gradient method, which is implemented using the automatic differentiation framework PyTorch [24] and the Adam optimizer [16]. The joint design and estimation method is summarized in Algorithm 1 and the code is made available here TO APPEAR.
Input: $t$,$(\check{\beta}^{b},\check{C}_{\beta^{b}})$, $(\widehat{v}^{b-1},C_{v^{b-1}})$
  if $t\text{ mod }B=0$ then
     estimate $\widehat{\beta}^{b},C_{\beta^{b}},\widehat{v}^{b}$ using (20)
     predict $\check{\beta}^{b+1},\check{C}_{\beta^{b+1}}$ from (19) using $u^{b}$,$\widehat{\beta}^{b}$,$C_{\beta^{b}}$
     $(\check{\beta}_{t+1},\check{C}_{\beta_{t+1}}):=(\check{\beta}^{b+1},\check{C}_{\beta^{b+1}})$
  else
     predict $\check{\beta}_{t+1},\check{C}_{\beta_{t+1}}$ from (19) using $u_{t}$,$\check{\beta}_{t}$, $\check{C}_{\beta_{t}}$
  end if
  form matrix (17) using $(\check{\beta}_{t+1},\check{C}_{\beta_{t+1}})$
  form cost in (10) using $C_{t}$
  minimize (10) using first-order method and apply (24)
  return $u_{t+1}^{*}$ and $\widehat{\theta}_{t+1}$
Algorithm 1 Adaptive Input Design and Estimation
4 Numerical Experiments
The suitable choice of block size $B$ in the block-wise estimator depends on the nature of the nonlinear model; ideally, we select the smallest $B\geq d_{\theta}+d_{x}+d_{y}$ that achieves good initial convergence. Also the design horizon $k$ is a choice that must balance computational cost and performance gains. In our examples, we found that $k>3$ is preferred. For all cases, we set the gain in (10) to $\gamma=400\gg d_{\theta}$.
The experiments below are initialized using $B$ samples with prbs inputs. The initial $\check{\beta}_{0}$ is randomized using a Gaussian with zero mean, and set to be uninformative using a large error covariance matrix $\check{C}_{\beta^{0}}=10^{4}\cdot I$. We assume accurate initial estimates $\widehat{v}_{0}$, setting $C_{v_{0}}$ such that there is a 0.1% error in the noise standard deviations. For the first-order method solving (10), we use a stepsize of $10^{-3}$.
To provide a succinct evaluation of the parameter estimates, we use the normalized mean squared error
$\text{MSE}_{t}=\sum^{d_{\theta}}_{k=1}\frac{\mathbb{E}[|\theta_{k}-\widehat{\theta}_{k,t}|^{2}]}{\theta^{2}_{k}}$
and the corresponding (approximate) Cramér-Rao bounds obtained by evaluating the fim in (17) at the unknown parameters.
4.1 Pendulum system
[…]
An example of the designed inputs is shown in Figure 1a. The adaptive design method starts by exceeding the angular boundaries, but as information about the parameters and states increases it is capable of planning ahead and rapidly steers the system within the constraint set.
4.2 Cruise control system
Next, we consider the following velocity model of a vehicle:
$x_{t+1}=x_{t}+\big{(}\theta_{1}x^{2}_{t}\text{sign}(x_{t})+\theta_{2}u_{t}\big{)}\Delta{T},$ (25)
[…]
In Figure 2, we see that the adaptively designed input increases the velocity until it reaches the limit and then returns the vehicle to stand still, achieving significantly more informative experiments than the unconstrained prbs. Moreover, further reductions of the MSE are possible when the gain $\gamma=0$, thus illustrating the trade off between informativeness and adherence to state constraints. On the other hand, using the popular A-criterion yields a notably different velocity that results in larger estimation errors.
Figure 2: Results for cruise control system (25). (a) Example of resulting velocity $x_{t}$ using adaptive designs with and without ()(*)( ∗ ) state constraints. (b) Squared estimation errors of estimator along with CRB. 100 Monte Carlo simulations.
4.3 Two tanks in series
prbs is near optimal for a wide class of linear systems and we are therefore interested in comparing its performance with that of the adaptive design when the system dynamics are nearly linear. Specifically, we consider the two-tank system:
$\begin{split}x_{t+1}^{(1)}&=x_{t}^{(1)}+\frac{(\theta_{1}u_{t}-\theta_{2}\sqrt{x_{t}^{(1)}})}{A_{1}}\Delta{T}\\ x_{t+1}^{(2)}&=x_{t}^{(2)}+\frac{(\theta_{2}\sqrt{x_{t}^{(1)}}-\theta_{3}\sqrt{x_{t}^{(2)}})}{A_{2}}\Delta{T},\end{split}$ (26)
where $x^{(1)}$ and $x^{(2)}$ represent the water levels in each tank and the known cross-sectional areas are denoted $A_{1}$ and $A_{2}$. The inlet flow rate to the first tank is determined by $u$, while the outlet flow rates are influenced by the square root of the respective water levels. The vector $\theta$ contains the unknown coefficients of the inlet, connecting, and outlet valves. Here $H=I$ in the measurement model (3). In the experiment, we set $\theta=\begin{bmatrix}1&0.2&0.15\end{bmatrix}^{\top}$, noise level $\sqrt{v}=0.01$ m, and sampling period $\Delta T=0.2$ s. The input is constrained as $u_{t}\in[0,0.5]$ and we want to prevent overflow of either tanks, which have a maximum level $x_{\max}=3$. The system is initialized at levels $x_{0}=\begin{bmatrix}1&1\end{bmatrix}^{\top}$. For the adaptive design, we set $k=4$ and $B=6$.
Figure 3 shows an example realization of the input design which indeed is close to a binary signal, but unlike prbs, it ensures that there is no tank overflow. The simulations show that the adaptive design tends to cyclically empty the first tank. Figure 4 confirms that the performance of the designed input for this nearly linear system is similar to that of the unconstrained prbs after about 20 samples. This is true even when $\gamma=0$.
Figure 3: Example for two-tank system (26). (a) prbs and adaptively designed input signal. (b) Corresponding water levels $x_{t}^{(1)}$ and $x_{t}^{(2)}$.
Figure 4: Squared estimation errors of parameter estimator along with CRB for two-tank system (26). ()(*)( ∗ ) denotes no active state constraints. 100 Monte Carlo simulations.
5 Conclusion
We have developed a joint input design and parameter estimation method for identifying nonlinear system models through the sequential acquisition of measurements. This is based on a receding horizon framework that utilizes a new adaptive, scale-invariant design criterion and a standarized penalization of the state constraint violation. The framework utilizes a sequential estimator based on the maximum likelihood principle, which was evaluated statistically.
The new methodology can lead to substantial gains in estimation accuracy and, therefore, reduced experiment time while steering the system within the operational constraint set.
{ack}