### v2

# Adaptive Input Design for Nonlinear System Identification with Operational Constraints

Abstract
We consider the problem of joint input design and parameter estimation for identifying nonlinear system models through the sequential acquisition of measurements while adhering to system constraints. We utilize a receding horizon approach and propose a new adaptive input design criterion, which is tailored to continuously updated parameter estimates, along with a new sequential parameter estimator. We demonstrate the ability of the method to design informative experiments online, while steering the system within operational constraints.
I Introduction
We consider discrete-time nonlinear dynamical systems of the form
$x_{t+1}=f_{\theta}(x_{t},u_{t}),$ (1)
where both states $x_{t}$ and parameters $\theta$ are unknown. Using the inputs $u_{t}$ and measurements of the system outputs $y_{t}$, we wish to estimate $\theta$. The goal is to choose $u_{t}$ so as to make the data informative about $\theta$, while respecting the system’s operating constraints. This paper thus tackles the challenging joint problem of adaptive input design and parameter estimation for nonlinear systems.
Input design for linear systems has been researched at least since the 1960s. A classic choice for $u_{t}$ is a pseudorandom binary sequence (prbs) [10] due to its excitation properties and ease of implementation. More efficient input designs are possible by leveraging a known structure of the system using the statistically motivated Fisher information matrix, $J$, or asymptotic error covariance matrices, in the design criterion [17, 7, 22, 12, 30]. For tractability, many approaches design the input signals in the frequency domain and then synthesize $u_{t}$ in the time domain [4, 13, 6, 1]. The high cost of conducting identification experiments motivates the development of efficient input design schemes that continuously adapt $u_{t}$ to the information available up to time $t$. This implies a closed-loop method for joint design and identification. It was shown in [2, 16] that for finite-impulse response and autoregressive system models, optimal adaptive designs under amplitude constraints correspond to binary sequences [12]. More recently, an adaptive input design for output-error models was considered in [32]. Informative experiments may, however, have adverse effects on the closed-loop system behaviour. The balance between sufficiently perturbing input signals and control performance was initially addressed by Feldbaum [8] and is a trade off considered in recent work on nonlinear system identification [27, 3].
The focus of this work is on the adaptive design of input signals $u_{t}$ that is informative about $\theta$ while steering the system states $x_{t}$ within operational constraints. The joint design and estimation problem is particularly challenging in the case of nonlinear systems, not least because standard one-step recursive estimation methods, such as the Extended Kalman filter, are highly sensitive to initial parameter and state estimates. Goodwin [11] introduced an open-loop input design method for nonlinear systems that penalizes designs in which the estimated states violate the constraints. This is achieved by linearizing the state-space model and designing $u_{t}$ over a fixed horizon. However, the method requires a good prior estimate of $\theta$ and the use of $\text{tr}\{J^{-1}\}$ mix estimation errors in each element of $\theta$ that typically have incommensurable units. More recently, Flila et al. [9] proposed the use of a receding-horizon approach to steer the nonlinear system within output constraints. This was achieved using a penalization technique and a design criterion that approximates $1/\text{tr}\{J\}$. The method was demonstrated for a thermal process model with temperature constraints. While computationally convenient, this design criterion is neither scale invariant nor does it lower bound the resulting estimation errors. Moreover, the statistical efficiency of the method is not studied.
In this paper, we aim to address the above challenges of adaptive designs for nonlinear system identification. Specifically, the principal contributions are:
  • a new adaptive design criterion tailored to the setting in which the parameter estimates are continuously updated;
  • an adaptive input design method with efficient parameter estimation that steers the system within operational constraints while achieving accurate estimates and thus reducing experiment time.
In the subsequent section, the problem of interest is formalized, followed by the introduction of an adaptive and scale-invariant design criterion. The design problem is rendered tractable using a bound on the criterion and relaxing the state constraints. The method description is concluded by proposing a joint parameter and state estimation method suitable for the problem. Statistical evaluations are presented using two different nonlinear systems as examples, after which we conclude the paper.
II Problem Formulation
The system output is assumed to follow the measurement model
$y_{t}=Hx_{t}+e_{t}$ (2)
where $y_{t}$ is a $d_{y}\times 1$ vector and $e_{t}\sim\mathcal{N}(0,V)$ is a white noise process with a diagonal covariance matrix $V$ with diagonal elements given by $v$. Sensor specifications typically provide relatively accurate estimates of the variances $v$. The dimensions of $u_{t}$, $\theta$ and $x_{t}$ are $d_{u}\times 1$, $d_{\theta}\times 1$ and $d_{x}\times 1$, respectively.
A sequence of input-output signals $\{(u_{0},y_{1}),\dots,(u_{t-1},y_{t})\}$ provides information about $\theta$ in (1) via (2). At any $t$, our task is to design subsequent inputs $u_{t+i}$ so that the resulting observations provide the maximal additional information about $\theta$, while steering the system within user-defined operational constraints:
$x_{\min}\leq x_{t+i}\leq x_{\max},$ (3)
which can be time varying. This is readily extended to constraints on the outputs $Hx_{t}$ via (2). A fundamental challenge here is that $x_{t+i}$ is given by the forward recursion of (1):
$\begin{split}x_{t+i}&=f_{\theta}(f_{\theta}(\dots f_{\theta}(x_{t},u_{t})\dots u_{t+i-2}),u_{t+i-1})\\ &\equiv f^{i}_{\theta}(x_{t},u_{t:t+i-1}),\;\forall i>0\end{split}$ (4)
where $\theta$ and $x_{t+1}$ have to be estimated from data. Thus (4) renders the constraints (3) nonlinear in the designed inputs, which must also satisfy
$u_{\min}\leq u_{t+i}\leq u_{\max}.$ (5)
Let $\mathcal{U}$ denote the set of input sequences that satisfy (5) and let $\mathcal{X}$ be the set of resulting states that satisfy (3). It is assumed that a feasible sequence exists in $\mathcal{U}$. To tackle the computational challenges ensuing from (4), we will leverage advances in automatic differentiation methods.
[…]
III Method
Suppose a $k$-step input sequence
$U_{t}=[u_{t},u_{t+1},\dots,u_{t+k-1}]$ (6)
is to be designed. It yields corresponding outputs $\{y_{t+1},\dots,y_{t+k}\}$. We begin by formulating a suitable adaptive design criterion for $U_{t}$ that quantifies the additional information gained about $\theta$ relative to what was available at $t$.
III-A Adaptive design criterion
The likelihood of the $k$ new outputs depends on $\theta$, $v$ and the unknown state $x_{t}$ via (4). Implementing $U_{t}$ yields the following Fisher information matrix (fim) of the unknown variables:
$J_{\theta,x_{t},v}(U_{t})=\mathbb{E}\left[\begin{bmatrix}\frac{\partial\ell}{\partial\theta}\\ \frac{\partial\ell}{\partial x_{t}}\\ \frac{\partial\ell}{\partial v}\\ \end{bmatrix}\begin{bmatrix}\frac{\partial\ell}{\partial\theta}\\ \frac{\partial\ell}{\partial x_{t}}\\ \frac{\partial\ell}{\partial v}\\ \end{bmatrix}^{\top}\right],$ (7)
where $\ell$ denotes the log-likelihood function of the data. It follows from the Cramér-Rao inequality [28, 21] that
$\begin{split}C_{t+k}=\begin{bmatrix}I_{d_{\theta}}&0\end{bmatrix}J^{-1}_{\theta,x_{t},v}\end{split}\begin{bmatrix}I_{d_{\theta}}&0\end{bmatrix}^{\top}$ (8)
is the minimum achievable error covariance matrix for any unbiased estimator $\widehat{\theta}_{t+k}$ that incorporates the $k$ new samples.
Let $\widehat{\theta}_{t}$ denote a prior estimate at time $t$ with error covariance matrix $C_{t}\succeq C_{t+k}$. Then
$\boxed{\mathcal{J}(U_{t};\theta,x_{t},v)=\text{tr}\left\{C_{t+k}C^{-1}_{t}\right\}\leq d_{\theta},}$ (9)
quantifies the relative error improvement gained by the design $U_{t+1}$. The upper bound $d_{\theta}$ is attained in the extreme case of uninformative experiments. The criterion (9) bounds the mean squared-error of $\widehat{\theta}_{t+k}$, standardized to the errors at $t$:
$\mathcal{J}\leq\mathbb{E}\left[\|\theta-\widehat{\theta}_{t+k}\|^{2}_{C^{-1}_{t}}\right],$ (10)
via the Cramér-Rao inequality. Thus $\mathcal{J}$ is an L-criterion [7, ch. 2.9] adapted to the information about $\theta$ available at $t$ that renders the errors in the components of $\theta$ commensurable.
A receding horizon formulation for adaptive input design is then
$\begin{split}U^{*}_{t}&=\operatorname*{arg\,min}_{U_{t}\in\mathcal{U}}\mathcal{J}(U_{t};\theta,x_{t},v)\\ &\;\text{subject to }x_{t+i}\in\mathcal{X},\forall i=1,\dots,k,\end{split}$ (11)
which is evaluated at estimates $\widehat{\theta}$, $\widehat{x}_{t}$ and $\widehat{v}$ given in Section III-C. However, both the constraints in (11) and the log-likelihood for all data render the problem rather intractable. Next, we address both issues.
III-B Relaxation
The Gaussian likelihood function for $y_{1:t}$ can be parameterized using an initial state $x_{0}$ and does not depend on $U_{t}$. To obtain a tractable fim using all data, consider summarizing the past data $y_{1:t}$ using maximum likelihood estimates (mle) $\widehat{\theta},\widehat{x}_{t}$ and $\widehat{v}$, whose distribution is denoted $q(\widehat{\theta},\widehat{x}_{t},\widehat{v}|\theta,x_{t},v)$. Under certain regularity conditions, $q$ is asymptotically normal and the estimator efficient so that with a Gaussian data model, its asymptotic covariance matrix is block-diagonal:
$\mathbb{V}\left[\begin{bmatrix}\widehat{\theta}\\ \widehat{x}_{t}\\ \widehat{v}\end{bmatrix}\right]=\begin{bmatrix}P&0\\ 0&Q\end{bmatrix},$ (12)
where the lower block $Q$ belongs to $\widehat{v}$ which is thus uncorrelated with $\widehat{\theta}$ and $\widehat{x}_{t}$ [31, appendix B]. We will therefore take $q$ to be a Gaussian distribution with a given covariance matrix (12) as summary of the past data. This is most conservative distribution in the sense that it yields the smallest fim [29, 25], and, moreover, corresponds to a factorization of the past likelihood of $y_{1:t}$ under certain regularity conditions for the asymptotic sufficiency of mle [15].
The corresponding likelihood function then becomes
~$\begin{split}\widetilde{\ell}&=\ln q(\widehat{\theta},\widehat{x}_{t},\widehat{v}|\theta,x_{t},v)+\sum^{k}_{i=1}p(y_{t+i}|u_{t:t+i-1},\theta,x_{t},v)\\ &=-\frac{1}{2}\left\|\begin{bmatrix}\theta\\ x_{t}\\ \end{bmatrix}-\begin{bmatrix}\widehat{\theta}\\ \widehat{x}_{t}\\ \end{bmatrix}\right\|_{P^{-1}}^{2}-\frac{1}{2}\left\|v-\widehat{v}\right\|_{Q^{-1}}^{2}\\ &\quad-\frac{1}{2}\sum_{i=1}^{k}\|\varepsilon_{t+i}\|_{V^{-1}}^{2}-\frac{k}{2}\ln{|V|}+\text{const.},\end{split}$ (13)
where we used the prediction error $\varepsilon_{t+i}\equiv y_{t+i}-Hf^{i}_{\theta}(x_{t},u_{t:t+i-1})$ for compactness. The corresponding conservative fim (7) becomes [28]:
$\begin{split}\widetilde{J}_{x_{t},\theta,v}&=\begin{bmatrix}\widetilde{J}_{\theta}&\widetilde{J}_{\theta x_{t}}&0\\ \widetilde{J}_{\theta x_{t}}^{\top}&\widetilde{J}_{x_{t}}&0\\ 0&0&\widetilde{J}_{v}\end{bmatrix}\\ &=\begin{bmatrix}P^{-1}+\sum_{i=1}^{k}E^{\top}_{t+i}V^{-1}E_{t+i}&{0}\\ {0}&Q^{-1}+\frac{k}{2}V^{-2}\end{bmatrix},\end{split}$ (14)
where
$E_{t+i}=H\begin{bmatrix}\frac{\partial f^{i}_{\theta}(x_{t},u_{t:t+i-1})}{\partial\theta}&\frac{\partial f^{i}_{\theta}(x_{t},u_{t:t+i-1})}{\partial x_{t}}\end{bmatrix}$
is the Jacobian of the prediction error $\varepsilon_{t+i}$. Applying (14) to (8), we finally obtain
$\widetilde{C}_{t+k}\equiv(\widetilde{J}_{\theta}-\widetilde{J}_{\theta x_{t}}\widetilde{J}_{x_{t}}^{-1}\widetilde{J}_{\theta x_{t}}^{\top})^{-1}\succeq C_{t+k}.$ (15)
Inserting (15) into the criterion (9) yields a tractable, conservative minimum error bound $\widetilde{\mathcal{J}}\geq\mathcal{J}$, which becomes tight under asymptotic sufficiency of the mle.
We now turn to relaxing the state constraints $\mathcal{X}$ in (11). We aim for designs that continually steer the state to satisfy the operational constraints. This is achieved by introducing a penalty term to the design problem (11). Whereas $\mathcal{J}$ measures a standarized squared deviation of $\widehat{\theta}_{t+k}$ from the unknown $\theta$, we let penalty $\mathcal{J}_{\mathcal{X}}$ measure a standardized squared exceedance of $x_{t+i}$ from the state boundaries. Specifically,
$\begin{split}&\mathcal{J}_{\mathcal{X}}(U_{t};\theta,x_{t})=\frac{1}{k}\times\\ &\sum_{i=1}^{k}\|(Sx_{\min}-Sx_{t+i})^{+}+(Sx_{t+i}-Sx_{\text{max}})^{+})\|^{2},\end{split}$ (16)
where $S$ is a diagonal matrix that standardizes violations from the constraints. That is, $[S]_{i,i}=1/(x^{(i)}_{\max}-x^{(i)}_{\min})$ or 0, depending on whether $x_{i}$ is constrained. The relaxed form of (11) then becomes
$U^{*}_{t}=\operatorname*{arg\,min}_{U_{t}\in\mathcal{U}}\widetilde{\mathcal{J}}(U_{t};\theta,x_{t},v)+\gamma\mathcal{J}_{\mathcal{X}}(U_{t};\theta,x_{t}),$ (17)
where $\gamma\geq 0$ is a weight that controls the adherence to the state constraints. Note that when the state sequence $x_{t+i}$ is in $\mathcal{X}$, the cost $\mathcal{J}_{\mathcal{X}}$ vanishes. By setting, for instance, $\gamma>d_{\theta}/(0.01)^{2}$ the cost of exceeding the state boundaries by 1%, exceeds the cost of using uninformative inputs. This incentivizes designs that steer the system within the state constraint set $\mathcal{X}$.
III-C Online estimator
Joint estimation of $(\theta,x_{t},v)$ for the design (17) involves $d_{\theta}+d_{x}+d_{y}$ parameters and is a challenging task for standard recursive estimators. To reduce the sensitivity to poor initial estimates, we update $\widehat{\theta}$ periodically using blocks of $b\geq d_{\theta}+d_{x}+d_{y}$ samples. At the same time, we maintain a running estimation of $x_{t}$ using the model in (1).
Suppose $t$ is a multiple of $b$, so that $\tau=t-b$ indicates the end of the previous block of samples. Let $(\widehat{\theta}_{\tau},\widehat{x}_{\tau|\tau},\widehat{v}_{\tau})$ denote the prior estimates at $\tau$. Using the likelihood function along with the $b$ subsequent samples, new estimates $\widehat{\theta}_{t}$, $\widehat{v}_{t}$, and a smoothed estimate $\widehat{x}_{\tau|t}$ is obtained by solving:
$\begin{split}&(\widehat{\theta}_{t},\widehat{x}_{\tau|t},\widehat{v}_{t})=\operatorname*{arg\,min}_{\theta,x_{\tau},v}\;\sum_{i=1}^{b}\|\varepsilon_{\tau+i}\|_{V^{-1}}^{2}+b\ln{|V|}\\ &+\left\|\begin{bmatrix}\theta\\ x_{\tau}\end{bmatrix}-\begin{bmatrix}\widehat{\theta}_{\tau}\\ \widehat{x}_{\tau|\tau}\end{bmatrix}\right\|_{P_{\tau|\tau}^{-1}}^{2}+\|v-\widehat{v}_{\tau}\|^{2}_{Q^{-1}_{\tau}},\end{split}$ (18)
which is a maximum likelihood estimator. The covariance matrices are therefore approximated using the corresponding inverse fim:
$P_{\tau|t}=\begin{bmatrix}\widetilde{J}_{\theta}&\widetilde{J}_{\theta x_{\tau}}\\ \widetilde{J}_{\theta x_{\tau}}^{\top}&\widetilde{J}_{x_{\tau}}\end{bmatrix}^{-1}\quad\text{and}\quad Q_{t}=\widetilde{J}_{v}^{-1}$ (19)
evaluating (14) with plug-in estimates. The state estimate in (18) and the joint covariance (19) are then sequentially updated to $t$:
$(\widehat{x}_{\tau|t},P_{\tau|t})\rightarrow(\widehat{x}_{\tau+1|t},P_{\tau+1|t})\rightarrow\dots\rightarrow(\widehat{x}_{t|t},P_{t|t})$
using the prediction step of the unscented Kalman filter [14, 23, 5].
The procedure is described as follows. Let $\widehat{z}_{t}=\begin{bmatrix}\widehat{\theta}\\ \widehat{x}_{t}\end{bmatrix}$ denote an initial estimate with an error covariance matrix $P_{t}$. The unscented transform is a deterministic sampling method that produces a one-step prediction, and its approximate error covariance, using $S=2d_{z}+1$ samples with $d_{z}=d_{\theta}+d_{x}$, is given by:
$\begin{split}(\widehat{z}_{t+1},P_{t+1})&=T(\widehat{z}_{t},P_{t};u_{t}).\end{split}$ (20)
The augmented state estimate is computed as
$\widehat{z}_{t+1}=\frac{1}{S}\sum_{s=0}^{S-1}w^{(s)}z^{(s)}_{t+1},\text{ where}z^{(s)}_{t+1}=\begin{bmatrix}\widehat{\theta}^{(s)}\\ f_{\widehat{\theta}^{(s)}}(\widehat{x}^{s}_{t},u_{t})\end{bmatrix}$
and
$z_{t}^{(s)}=\begin{cases}\widehat{z}_{t}&\text{for}\quad s=0\\ \widehat{z}_{t}+\eta^{1/2}[P_{t}^{1/2}]_{s}&\text{for}\quad s=1,\dots,d_{z}\\ \widehat{z}_{t}-\eta^{1/2}[P_{t}^{1/2}]_{s}&\text{for}\quad s=d_{z}+1,\dots,2d_{z}\end{cases}.$ (21)
The weights $w^{(s)}$ and $\eta$ are given in [14, 23, 5]. The covariance matrix $P_{t}$ is computed similarly via outer products of the samples. Eq. (20) is applied sequentially to produce running predictions of $x_{t}$ for use in (17).
III-D Adaptive design method
Using the joint estimate $(\widehat{\theta},\widehat{x}_{t},\widehat{v})$ described above, we can finally evaluate the criterion in (17). This design problem can be converted into an unconstrained form by transforming the inputs $\mathcal{U}\rightarrow\mathbb{R}^{d_{u}}$ as follows:
$\widetilde{U}_{t}=[\widetilde{u}(u_{t}),\dots,\widetilde{u}(u_{t+k-1})]$ (22)
where
$\widetilde{u}(u_{t})=g\big{(}(u_{t}-u_{\text{min}})\oslash(u_{\text{max}}-u_{\text{min}})\big{)}$ (23)
and where $g$ is a differentiable and invertible function, e.g., the element-wise inverse sigmoid function $[g(z)]_{i}=\ln{\frac{z_{i}}{1-z_{i}}}$.
Thus (22) is unconstrained and computationally convenient for adaptive input design. The map $\widetilde{U}_{t}$ recovers the original input space by:
$U_{t}=[\widetilde{u}^{-1}(\widetilde{u}_{t}),\dots,\widetilde{u}^{-1}(\widetilde{u}_{t+k-1})]$ (24)
where $\widetilde{u}^{-1}(\tilde{u}_{t})=g^{-1}(\tilde{u}_{t})\odot(u_{\max}-u_{\min})+u_{\min}$ and $[g^{-1}(z)]_{i}=\frac{1}{1+e^{-z_{i}}}$. Using the transformed inputs $\widetilde{U}_{t}$, by substituting (24) in (17), results in an unconstrained optimization problem. At $t$, the input
$u^{*}_{t}=\widetilde{u}^{-1}(u^{*}_{t})$
is computed and applied in the subsequent time step.
The adaptive input design and online estimation method can now be summarized as follows:
  1. 1.
    Observe $b$ new samples, solve (18) and compute (19).
  2. 2.
    Update estimate of $x_{t}$ and $P_{t}$ using (20).
  3. 3.
    Solve (17) in unconstrained form and apply input $u^{*}_{t}$.
  4. 4.
    Update $t:=t+1$
  5. 5.
    Go to 1) if $t$ is a multiple of $b$, otherwise go to 2).
To solve the non-convex minimization problem in (18), we employ a numerical optimization strategy suitable for rugged cost surfaces. Specifically, we utilize the Basin Hopping algorithm [19, 24] to locate an approximate solution to (18), then using a quasi-Newton limited-memory BFGS algorithm [20] for local optimization. To solve the design problem (17), we leverage an available automatic differentiation method [26] alongside the Adam optimizer [18], using a constant stepsize of $10^{-3}$. The code implementations are available at TODO: URL. The runtime of the design method using a standard laptop with single thread CPU is in the range of 0.5 to 5 seconds, depending on $k$ and model complexity.
IV Numerical Experiments
The experiments presented below are initialized using a block of $b$ samples generated from prbs inputs. The initial prior estimates, $\widehat{\theta}_{0}$ and $\widehat{x}_{1|0}$, are randomly drawn from a Gaussian distribution with zero mean and chose to be uninformative by assigning a large covariance matrix $P_{1|0}=10^{4}\cdot I$. The initial noise estimates $\widehat{v}_{0}$ are presumed to be accurate, and their covariance matrix $Q_{0}$ is configured to correspond to a 0.1% uncertainty in the noise standard deviations. For all the experiments, the weight in (17) was set to $\gamma=400\gg d_{\theta}$, where $d_{\theta}$ equals 2 or 3 below.
To provide a succinct evaluation of the parameter estimates, we use the normalized mean squared error
$\textsc{mse}_{t}=\sum^{d_{\theta}}_{k=1}\frac{\mathbb{E}[|\theta_{k}-\widehat{\theta}_{k,t}|^{2}]}{\theta^{2}_{k}}$
and the corresponding (approximate) Cramér-Rao bounds (crb) obtained by evaluating the fim in (15) at the true parameters. In addition, we quantify the (relative) violation of operational constraints (3) using:
$\textsc{ocv}_{t}=\mathbb{E}\left[\|(Sx_{\min}-Sx_{t})^{+}+(Sx_{t}-Sx_{\text{max}})^{+})\|\right],$
with $S$ from (16).
IV-A Pendulum system
Figure 1: Example realization of pendulum system. Angle $x_{t}^{(1)}$ when applying two different prbs inputs and the adaptive design that steers the angle within the interval $[-45^{\circ},45^{\circ}]$.
Figure 2: Pendulum system (25) with input $|u_{t}|\leq 10$ starting at sample $t=8$. (a) Relative violation of operational constraint (3). (b) Mean-square errors of parameter estimator along with (approximate) Cramér-Rao bounds. The estimation errors for prbs2 were too large to be included in the plot. Results are based on 100 Monte Carlo runs.
First, we consider a nonlinear pendulum model:
$\begin{split}x_{t+1}^{(1)}&=x_{t}^{(1)}+x_{t}^{(2)}\Delta{T}\\ x_{t+1}^{(2)}&=x_{t}^{(2)}+(\theta_{1}\sin{x_{t}^{(1)}}+\theta_{2}u_{t})\Delta{T},\end{split}$ (25)
[…]
For comparison, we use two prbs designs. The first one, prbs1, only satisfies input amplitude constraints, whereas the second one, prbs2, is tuned to the best of our ability to satisfy the operational constraints during an experiment. An example realization is provided in Fig. 1, where we see that the adaptive design quickly steers the unknown system within the constraints. In contrast, prbs2 yields an uninformative experiment, while prbs1 is informative but grossly violates the constraints. A more systematic evaluation is provided in Fig. 2, where we see that the adaptive design also yields a lower crb and estimation error than prbs1, which is not subject to any state constraints.
IV-B Two tanks in series
prbs is near optimal for a wide class of linear systems and we are therefore interested in comparing its performance with that of the adaptive design when the system dynamics are nearly linear. Specifically, we consider the two-tank system, where the output of one tank is the input to the second:
$\begin{split}x_{t+1}^{(1)}&=x_{t}^{(1)}+\frac{(\theta_{1}u_{t}-\theta_{2}\sqrt{x_{t}^{(1)}})}{A_{1}}\Delta{T}\\ x_{t+1}^{(2)}&=x_{t}^{(2)}+\frac{(\theta_{2}\sqrt{x_{t}^{(1)}}-\theta_{3}\sqrt{x_{t}^{(2)}})}{A_{2}}\Delta{T},\end{split}$ (26)
where $x^{(1)}$ and $x^{(2)}$ represent the water levels in each tank and the known cross-sectional areas are denoted $A_{1}$ and $A_{2}$. The inlet flow rate to the first tank is determined by $u_{t}$. The outlet flow rates are influenced by the square root of the respective water levels. The vector $\theta$ contains the unknown coefficients of the inlet, connecting, and outlet valves. Here $H=I$ in the measurement model (2). In the experiment, we set $\theta=\begin{bmatrix}1&0.2&0.15\end{bmatrix}^{\top}$, noise level $\sqrt{v}=0.01$ m, and sampling period $\Delta T=0.2$ s. The input is constrained as $u_{t}\in[0,0.5]$ and we want to prevent overflow of either tanks, which have a maximum level $x_{\max}=3$. The system is initialized at levels $x_{0}=\begin{bmatrix}1&1\end{bmatrix}^{\top}$. For the adaptive design, we set $k=4$ and $b=6$.
Figure 3 shows an example realization. The adaptive design yields a binary signal as anticipated, but unlike prbs, it prevents tank overflow. The simulations show that the adaptive design tends to cyclically empty the first tank. In addition to these designs, we also include an unconstrained adaptive design ($\gamma=0$) to the evaluation in Fig. 4. We see that while only the constrained design avoids overflow, it comes at a minor cost of estimation performance compared to the unconstrained design after about 20 samples.
Figure 3: Example realization of two-tank system (26), corresponding to prbs and adaptively designed input signal. The plot shows the water levels $x_{t}^{(1)}$ and $x_{t}^{(2)}$ for prbs inputs and the adaptive design that steers the levels within the interval $[0,3]$.
Figure 4: Results for two-tank system (26) (a) Relative violation of state constraints over time. (b) Standardized estimation errors of parameter estimator along with CRB. (*) denotes the design with no imposed state constraints (i.e., $\gamma=0$). Results are based on 100 Monte Carlo runs.
V Conclusion
We have developed an adaptive input design and sequential estimation method for identifying nonlinear system models. The method utilizes a new design criterion and a standarized penalization of the state constraint violation. Standard one-step joint state parameter estimators, such as the Extended Kalman filter, were too sensitive to poor initial estimates to provide useful information for the adaptive design procedure. We therefore developed and evaluated a more robust block-wise estimator based on the maximum likelihood principle, which was evaluated statistically. The computation time is near real-time on a standard laptop and future work can fruitfully explore more efficient implementations.
The new methodology can lead to substantial gains in estimation accuracy and, therefore, reduced experiment time while steering the nonlinear system within the operational constraint set.